# Dynamic Memory Compression: Retrofitting LLMs for Accelerated Inference

### v1

@@ 3.3 Practical Considerations
Implementing a variable-length cache without padding
DMC allows every head to learn a custom compression, which results in KV cache sequences with variable lengths across heads. One could implement this cache naïvely with padded tensors. Foreshadowing our results in Section 5.2, however, the largest efficiency gains during inference stem from the reduced memory footprint, which allows us to increase the batch size and dramatically improve throughput. In order to get these benefits, the memory cannot be wasted on storing KV cache as padded tensors.
Thus, we provide a simple custom implementation of attention in PyTorch, building on FlashAttention (Dao et al., 2022) and PagedAttention (Kwon et al., 2023), in which the heads can learn wildly different compression rates but which does not require padding. As an ablation, we also study a constrained variant (DMC-C) in which we force the heads in a given layer to maintain similar compression rates, which minimizes the padding necessary in naïve attention implementations; however, it significantly constrains the learned compression schema and leads to worse model quality.
@@
The calculation of partial accumulations, during training of DMC models Equation 9, for a sequence of $n$ tokens requires $O(n)$ sequential steps, therefore considerably slowing down the training. In order to improve the time complexity, we calculate Equation 9 at every time step $t$ independently over a short window of the last $w$ elements up to time $t$ (e.g., $w=12$). This enables us to reduce the span of computation to $O(w)$, provided that at least $n$ threads can execute the computation in parallel for each of the $n$ positions.
@@ 4 Experimental Setup
In our experiments, we evaluate strategies to retrofit a state-of-the-art Large Language Model (LLM), Llama 2 (Touvron et al., 2023), [^4] Obtained from https://huggingface.co/meta-llama. into a more efficient model across various sizes: 7B, 13B, and 70B. In addition to comparing the downstream performance of DMC with the original model, we also use Grouped Query Attention (GQA) as a main baseline, as it constitutes the most widespread strategy to ensure KV cache efficiency (Jiang et al., 2023).
@@
To equip the original Llama 2 with GQA, we perform the standard checkpoint conversion proposed by Ainslie et al. (2023): the key and value projection matrices are split by head. Then the resulting sub-matrices corresponding to heads in the same group are merged. Note that this results in a fixed CR during training.
As for DMC, we avoid the introduction of new parameters by re-purposing the first dimension from both the $\mathbf{k}_{t}$ and $\mathbf{q}_{t}$ representations, in order to predict segmentation decisions and importance scores. Setting $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ to zero triggers a significant increase in language modeling loss, by disrupting the attention scores. To counteract this spike in loss, we pre-train the model to disregard the first dimension of $\mathbf{k}_{t}$ and $\mathbf{q}_{t}$ in the attention calculations. Specifically, we load the raw pre-trained weights and up-train the model for 1 billion tokens (250 steps), annealing the values of $\mathbf{k}_{t}$ and $\mathbf{q}_{t}$ to 0 according to the following formula:
@@
where $t$ is the current step and $n_{t}=250$. After this initial phase, which allows the model to ignore the first dimension of keys and values for attention calculations, we start the main retrofitting phase. We set the constant from Equation 8 as $c=-5$ in order to perform no compression at the start.
@@
The volume of data for continued pre-training of DMC is contingent on the targeted KV cache compression ratio; a larger CR necessitates more data. We use a training schedule with 24B, 48B, and 72B tokens for training to 2×2$\times$2 ×, 3×3$\times$3 ×, and 4×4$\times$4 × compression, respectively. In Appendix D we include an ablation where we use a schedule with twice less data.
For DMC, we discovered that the annealing strategy was crucial. Starting the training without compression and its measured increase helps to preserve the original perplexity. Through extensive ablations (see Appendix D), we found that any significant increase of perplexity, even if recovered during continued pre-training, prevents the model from regaining its performance on downstream tasks. The target CR is linearly increased from 1×1$\times$1 × to 4×4$\times$4 × for the 7B and 13B models, and from 1×1$\times$1 × to 2×2$\times$2 × for the 70B model. [^5] For Llama 2 70B, we do not up-train to 4×4$\times$4 × because this LLM is already pre-trained with GQA 8×8$\times$8 ×..
Upon achieving the target compression ratio with the DMC model, we initiate a final solidifying phase wherein we: 1) continue up-training for an additional 8B tokens, 2) maintain a fixed compression ratio, and 3) implement a cosine learning rate schedule, annealing it down to 10% of the initial value. This phase aims at stabilizing the model with a specific, fixed compression ratio.
@@
Following Touvron et al. (2023), we evaluate models on a series of downstream tasks, including MMLU (Hendrycks et al., 2021) for factuality, HumanEval (Chen et al., 2021) for Python code generation, and several question-answering datasets for common-sense reasoning: PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), Arc-C and Arc-E (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and WinoGrande (Sakaguchi et al., 2021). We report the 5-shot performance on MMLU, average pass@1 scores for HumanEval, and average 0-shot performance on common-sense benchmarks (CS-QA). For pass@1 scores we use a temperature of 0.1 and nucleus sampling (Holtzman et al., 2019) with top-p $=$ 0.95.
@@ 5 Results
$*{}^{*}$ $\times^{*}$ 68.8 78.0 29.6
DMC $\times^{*}$ 68.8 77.9 29.9
Table 1: MMLU accuracy (Acc.), Commonsense Question Answering (CS-QA), Exact Match (EM), and Human-Eval Pass@1 for several scales (7B, 13B, and 70B) and compression rates (CRs; 1×1$\times$1 ×, 2×2$\times$2 ×, and 4×4$\times$4 ×) of Llama 2. (*) Unlike the 7B and 13B models, the 70B model was trained with GQA which compresses the KV cache 8×8$\times$8 ×. We apply additional 2×2$\times$2 × DMC compression during up-training to obtain a total compression of 16×16$\times$16 × .
@@
5.1 Main Results
We report the performance of the original LLM (equivalent to 1×1$\times$1 × CR) and efficient variants (DMC and GQA) in Table 1. For the original LLM we use results reproduced in our codebase as described in Appendix B.
(a) Inference throughput averaged over the generation of 1K tokens with 3K tokens of context (up to 4K in total). On the x-axis, we show the maximum batch size that fits in memory on a single GPU (7B and 13B) or two GPUs with tensor parallelism (70B) for the Vanilla, DMC 2$\times$, and DMC 4$\times$ models.
(b) Latency of next token generation. Solid lines denote measurements with the maximum batch size that fits on a single GPU. Dotted lines denote DMC 4$\times$ with the same batch size as the vanilla LLM.
@@ 5.1 Main Results
DMC vs Original First, comparing DMC with the original LLM, we note that it even increases the performance in MMLU and CS-QA at 2×2$\times$2 × CR for 7B and 13B and in Human-Eval for all model scales. We speculate that this is due to the additional up-training steps, which expose LLMs to new examples. For the other combinations of downstream tasks and scales at 4×4$\times$4 × CR, DMC-R incurs negligible degradation: -0.7 in MMLU and -0.3 in CS-QA for 7B, -0.3 in MMLU and CS-QA for 13B. This encouraging finding suggests that DMC is robust across different model scales even for 4×4$\times$4 × CR. Overall, the fact that DMC is in general close or superior to the original performance makes it suitable as a drop-in replacement for KV caching to achieve higher inference efficiency.
DMC vs GQA Moreover, Table 1 allows for comparing DMC with GQA for equivalent CRs (2×2$\times$2 × and 4×4$\times$4 ×). Overall, DMC surpasses GQA applied through up-training, in both CRs and in both scales (7B and 13B). For MMLU, the gap widens when we increase the CR. This holds true both at the smaller 7B scale (from +5.4 for 2×2$\times$2 × CR to +9.2 for 4×4$\times$4 × CR) and at the larger 13B scale (from +4.6 for 2×2$\times$2 × CR to +5.6 for 4×4$\times$4 × CR). For CS-QA and Human-Eval, on the other hand, we observe comparable gains over GQA across CRs and model scales. These findings illustrate that DMC should be preferred to GQA for retrofitting LLMs into variants that are more efficient at inference time.
70B: DMC and GQA However, many widely adopted LLMs were pre-trained with GQA which leads to the question of whether DMC and GQA can be used together to reap compounded benefits. To investigate this, we retrofit Llama 2 70B, which has been pre-trained with 8×8$\times$8 × GQA. We further compress its KV cache with DMC to 2×2$\times$2 × CR: this is equivalent to a cache 16×16$\times$16 × smaller than a vanilla LLM with neither GQA nor DMC. We observe that the performance remains unchanged, and conclude that DMC and GQA can be easily and successfully combined.
DMC vs DMC-C Finally, in Table 3 in Appendix E, we compare DMC with its Constrained variant DMC-C. In general, while remaining superior to GQA, DMC-C displays a significant degradation in several configurations, most notably 7B 4×4$\times$4 × where it records a drop of $-6.4$ in MMLU compared to the ceiling. On the other hand, DMC recovers all performance loss in DMC-C. When used in combination with custom attention implementations which does not require excessive padding, standard DMC should therefore be vastly preferred, as it retains the original LLM performance while reaping the advantages in memory efficiency fully.
Sample Efficiency To shed light on the sample efficiency of DMC and GQA, we report their performance on MMLU, CS-QA, and Codex-Eval across retrofitting steps in Figure 4. First, for a target CR of 2×2$\times$2 ×, it emerges how GQA cannot achieve the performance that DMC obtains after fine-tuning (at 8K steps) even after more than double the amount of fine-tuning (at 17K steps). This applies to both 7B and 13B scales. Figure 4 also reveals the importance of the fine-tuning phase in DMC: a limited amount of extra steps with a fixed CR recovers a significant amount of the original performance (especially for higher target CRs such as 4×4$\times$4 ×).
5.2 Throughput and Latency Measurements
To verify whether increased CRs result in concrete efficiency gains, in Figure 5 we present the performance properties of DMC, estimated within the NVIDIA Megatron-LM framework (Narayanan et al., 2021). Specifically, we run measurements on a single GPU (NVIDIA A100 80GB SXM or H100 SXM) in bfloat16 precision for Llama 7B and 13B. For Llama 70B, we run the same measurements on two GPUs of the same type with tensor parallelism. We feed the model with 2K tokens of English text, and generate additional 2K tokens in an auto-regressive manner to ensure that the model properly compresses its own generations. The reported throughput consists of the average over the last 1K tokens. We limit the sequence to 4K tokens to avoid issues with context length extrapolation, as this is the maximum length observed by Llama 2 during pre-training.
Figure 6: Heatmaps of average compression rates across layers (X-axis) and heads (Y-axis). Heads are arranged from the highest compression to the lowest top-down for clarity.
In order to maximize the utilization of the GPU, we increase the batch size to the maximum that fits into memory (see Appendix A for details). The compression of the KV cache with DMC allows for substantial increases in batch size and thus significant throughput gains. As shown in Figure 4(a), DMC 2×2$\times$2 × translates into an effective increase in tokens per second compared to the original LLM of >1.8×>1.8$\times$> 1.8 × for 7B, 13B, and 70B on both A100 and H100 GPUs. Similarly, DMC 4×4$\times$4 × translates between 3.4×3.4$\times$3.4 × and 3.7×3.7$\times$3.7 ×. This means that the efficiency boost observed in practice is very close to the theoretical limit. In addition, the extra memory saved with DMC could also be used to cache longer contexts.
@@ 5.2 Throughput and Latency Measurements
Finally, we compare the latency of the original LLM with DMC 4$\times$ in Figure 4(b). When we fit the largest possible batch size for either model, after generating approximately 2200 tokens, the inference time begins to scale linearly with the context size due to the increasingly dominating cost of reading the KV cache from the high bandwidth memory (HBM). On the other hand, if we choose to maintain the same batch size as for the original LLM also for DMC 4$\times$, the memory footprint of the KV cache is reduced and latency for longer contexts improves significantly.
@@
While we acknowledge that the behavior of LLMs at inference time depends on a multitude of factors and implementation details, our measurements in Figure 5 offer evidence that DMC increases throughput and reduces the latency of autoregressive generation with LLMs. We speculate that in the future, DMC might be used to grow the KV cache sub-linearly, which would provide an alternative between vanilla Transformers and State Space Models, where memory is constant (Fu et al., 2023; Gu & Dao, 2023).
5.3 Per-Head Learned Compression Rates

### v2

@@ 3.3 Practical Considerations
Storing a variable-length KV cache in memory
DMC allows every head to learn a custom compression, which results in KV cache sequences with variable lengths across heads. This poses difficulties in storing these sequences efficiently in an $n$-dimensional tensor, considering that they will be extended during auto-regressive generation by uneven amounts of tokens. However, such sequences can be easily stored in memory with little overhead using PagedAttention (Kwon et al., 2023), where new pages are allocated on-demand for every head separately. In Section 5.2 we present latency and throughput measured with an implementation based on FlashAttention (Dao et al., 2022) and PagedAttention.
@@
The calculation of partial accumulations, during training of DMC models Equation 10, for a sequence of $n$ tokens requires $O(n)$ sequential steps, therefore considerably slowing down the training. In order to improve the time complexity, we calculate Equation 10 at every time step $t$ independently over a short window of the last $w$ elements up to time $t$ (e.g., $w=12$). This enables us to reduce the span of computation to $O(w)$, provided that at least $n$ threads can execute the computation in parallel for each of the $n$ positions. During inference, this speed-up also applies to the prompt phase. However, the sliding window comes at a disadvantage as it needs to be cached during inference for heads that tend to accumulate more than $w$ tokens.
@@ 4 Experimental Setup
In our experiments, we evaluate strategies to retrofit a state-of-the-art Large Language Model (LLM), Llama 2 (Touvron et al., 2023)  [^4] Obtained from https://huggingface.co/meta-llama., into a more efficient model across various sizes: 7B, 13B, and 70B. In addition to comparing the downstream performance of DMC with the original model, we also use Grouped Query Attention (GQA) as a main baseline, as it constitutes the most widespread strategy to ensure KV cache efficiency (Jiang et al., 2023).
We also compare DMC with two KV cache eviction policies that do not require model retrofitting: Token Omission Via Attention (TOVA; Oren et al., 2024) and Heavy-Hitter Oracle (H2O; Zhang et al., 2023). H2O keeps in memory a fixed window of the most recent tokens, as well as additional heavy-hitter (H2) tokens. H2 tokens are chosen dynamically: specifically, they correspond to the tokens with the highest aggregated attention scores throughout the sequence. On the other hand, TOVA retains the top-k tokens based on the attention weights of the last token only. This means that for a given attention layer, at each decoding step, the token with the lowest attention score is omitted.
@@
To equip the original Llama 2 LLM with GQA, we perform the standard checkpoint conversion proposed by Ainslie et al. (2023): the key and value projection matrices are split by head. Then the resulting sub-matrices corresponding to heads in the same group are merged via averaging. Note that this results in a fixed compression during training.
As for DMC, we avoid the introduction of new parameters by re-purposing the first dimension from both the $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ representations, in order to predict segmentation decisions $\alpha_{t}$ and importance scores $\omega_{t}$. Setting $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ to zero triggers a significant increase in language modeling loss, by disrupting the attention scores. To counteract this spike in loss, we pre-train the model to disregard the first dimension of $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ in the attention calculations. Specifically, we load the pre-trained weights and up-train the model for 1 billion tokens (250 steps), annealing the values of $\mathbf{q}_{t}$ and $\mathbf{k}_{t}$ to 0 according to the following formula:
@@
where $t$ is the current step and $n_{t}=250$. After this initial phase, which allows the model to ignore the first dimension of keys and values for attention calculations, we start the main retrofitting phase, where the model learns to compress the KV representations.
@@
The volume of data for continued pre-training of DMC is contingent on the targeted KV cache compression ratio; a larger CR necessitates more data. We use a linear training schedule with 24B, 72B, 120B, and 168B tokens for training to 2×2$\times$2 ×, 4×4$\times$4 ×, 6×6$\times$6 ×, and 8×8$\times$8 × compression, respectively. In Appendix E we include an ablation where we use a schedule with twice less data.
We discovered that the right retrofitting strategy is crucial for DMC: starting the training without compression helps to preserve the original perplexity. Any significant increase in perplexity, even if recovered during continued pre-training, prevents the model from regaining its performance on downstream tasks (see the ablations in Appendix E). The target CR is linearly increased from 1×1$\times$1 × to 8×8$\times$8 × for the 7B and 13B models, and from 1×1$\times$1 × to 2×2$\times$2 × for the 70B model. [^5] For Llama 2 70B, we do not up-train to 4×4$\times$4 × because this LLM is already pre-trained with GQA 8×8$\times$8 ×.
Upon achieving the target compression ratio, we initiate a final solidifying phase wherein we: 1) continue up-training for an additional 8B tokens, 2) maintain a fixed compression ratio, and 3) implement a cosine learning rate schedule, annealing it down to 10% of the initial value. This phase aims at stabilizing the model with a specific, fixed compression ratio. Resource-wise, all retrofitting stages require roughly 8k, 16k, and 28k GPU hours of NVIDIA H100 for Llama 2 7B, 13B, and 70B respectively.
@@
Following Touvron et al. (2023), we evaluate models on a series of downstream tasks, including MMLU (Hendrycks et al., 2021) for factuality, HumanEval (Chen et al., 2021) for Python code generation, and several question-answering datasets for common-sense reasoning: PIQA (Bisk et al., 2020), BoolQ (Clark et al., 2019), Arc-C and Arc-E (Clark et al., 2018), HellaSwag (Zellers et al., 2019), and WinoGrande (Sakaguchi et al., 2020). We report the 5-shot performance on MMLU, average pass@1 scores for HumanEval, and average 0-shot performance on common-sense benchmarks (CS-QA). For pass@1 scores we use a temperature of 0.1 and nucleus sampling (Holtzman et al., 2020) with top-p $=$ 0.95.
We adapted TOVA and H2O to our codebase based on publicly released code. For a given CR, the total budget for both policies is calculated as $(1/\text{CR})\times n$ for MMLU and CS-QA tasks, where $n$ is the input length. For HumanEval, which instead involves generating more than one token, the initial budget is also $(1/\text{CR})\times n$ but then increases dynamically as we generate the answer. For H2O, the budget is split equally between the local window and heavy-hitter tokens.
@@ 5 Results
H2O 6$\times$ 45.9 54.9 2.4
TOVA 51.7 60.0 0.0
DMC 53.3 73.4 21.3
H2O 8$\times$ 41.8 50.9 1.8
TOVA 50.3 53.4 0.0
DMC 52.1 73.3 21.3
70B $\times^{*}$ 68.8 78.0 29.6
H2O $\times^{*}$ 68.7 74.1 18.3
TOVA 68.1 77.6 29.9
DMC 68.8 77.9 29.9
Table 1: MMLU accuracy, Commonsense Question Answering (CS-QA) accuracy averaged across 6 tasks, and HumanEval Pass@1 for several scales (7B, 13B, and 70B) and compression ratios (CRs; 1×1$\times$1 ×, 2×2$\times$2 ×, 4×4$\times$4 ×, 6×6$\times$6 ×, 8×8$\times$8 ×) of Llama 2. (*) Unlike the 7B and 13B models, the 70B model was trained with GQA which compresses the KV cache 8×8$\times$8 ×. We apply additional 2×2$\times$2 × DMC compression during retrofitting to obtain a total compression of 16×16$\times$16 ×.
5.1 Main Results
We report the performance of the original LLM (equivalent to 1×1$\times$1 × CR) and efficient variants (DMC, GQA, TOVA, and H2O) in Table 1. The results for the original LLM are those reproduced in our codebase as described in Appendix A.
DMC vs Original First, comparing DMC with the original LLM, we note that it even increases the performance in MMLU and CS-QA at 2×2$\times$2 × CR for 7B and 13B and in HumanEval for all model scales. We speculate that this is due to the additional up-training steps, which expose LLMs to new examples. For the other combinations of downstream tasks and scales at 4×4$\times$4 × CR, DMC incurs negligible degradation: -0.7 in MMLU and -0.3 in CS-QA for 7B, -0.3 in MMLU and CS-QA for 13B. At higher compression ratios of 6×6$\times$6 × and 8×8$\times$8 × CR DMC shows slight degradation but generally stays close to the original performance. For 7B, DMC exhibits -1.7 in MMLU at 6×6$\times$6 × CR, and -2.8 at 8×8$\times$8 × CR. For 13B performance drops are smaller than for 7B, DMC incurs -1.2 in MMLU and -0.1 in CS-QA at 6×6$\times$6 × CR, and -2.4 in MMLU and -0.2 in CS-QA at 8×8$\times$8 × CR. This encouraging finding suggests that DMC is robust across different model scales even for 8×8$\times$8 × CR. Overall, the fact that DMC is in general close or superior to the original performance makes it suitable as a drop-in replacement for vanilla KV caching to achieve higher inference efficiency.
DMC vs GQA Moreover, Table 1 allows for comparing DMC with GQA for equivalent CRs (2×2$\times$2 × and 4×4$\times$4 ×). Overall, DMC surpasses GQA applied post-hoc through up-training, in both CRs and in both scales (7B and 13B) where we compare them. For MMLU, the gap widens when we increase the CR. This holds true both at the smaller 7B scale (from +5.4 for 2×2$\times$2 × CR to +9.2 for 4×4$\times$4 × CR) and at the larger 13B scale (from +4.6 for 2×2$\times$2 × CR to +5.6 for 4×4$\times$4 × CR). For CS-QA and Human-Eval, on the other hand, we observe comparable gains over GQA across CRs and model scales. Notably, DMC with 8×8$\times$8 × compression outperforms GQA with 2×2$\times$2 × compression for both 7B and 13B model scales. These findings illustrate that DMC should be preferred to GQA for retrofitting LLMs into variants that are more efficient at inference time.
DMC vs Cache Eviction Policies Both H2O and TOVA evict tokens from memory according to the post-softmax attention scores, which makes them incompatible with efficient attention implementations like FlashAttention. In fact, this requires materializing the $n^{2}$-sized tensor of attention scores. As a result, while reducing the KV cache size, these policies may slow down inference. Moreover, comparing the performance of both cache eviction policies with DMC in Table 1, it emerges how they are almost comparable at CR 2$\times$ in MMLU and CS-QA; however, they drop significantly in accuracy with a higher CR of 4$\times$, 6×6$\times$6 ×, and 8×8$\times$8 ×. The gap is even more dramatic for HumanEval pass@1 at all CRs and scales, except only for 70B TOVA. At 6×6$\times$6 × CR, H2O and TOVA show significant drops in MMLU and CS-QA, with TOVA outperforming H2O. At 8×8$\times$8 × CR, the performance drops further for both policies, especially in HumanEval, where DMC consistently shows superior performance.
@@ 5.1 Main Results
70B: DMC and GQA Many widely adopted LLMs were pre-trained with GQA which leads to the question of whether DMC and GQA can be used together to reap compounded benefits. To investigate this, we retrofit Llama 2 70B, which has been pre-trained with 8×8$\times$8 × GQA. We further compress its KV cache with DMC to 2×2$\times$2 × CR: this is equivalent to a cache 16×16$\times$16 × smaller than a vanilla LLM with neither GQA nor DMC. We observe that the performance remains unchanged, and conclude that DMC and GQA can be easily and successfully combined.
Sample efficiency To shed light on the sample efficiency of DMC and GQA, we report their performance on MMLU, CS-QA, and HumanEval across retrofitting steps in Figure 4. First, for a target CR of 2×2$\times$2 ×, it emerges how GQA cannot achieve the performance that DMC obtains after fine-tuning (at 8K steps) even after more than double the amount of fine-tuning (at 17K steps). This applies to both 7B and 13B scales. Figure 4 also reveals the importance of the fine-tuning phase in DMC: a limited amount of extra steps with a fixed CR recovers a significant amount of the original performance (especially for higher target CRs such as 4×4$\times$4 ×).
5.2 Throughput and Latency Measurements
To verify whether increased CRs result in concrete efficiency gains, we present the performance properties of DMC in Figure 5, estimated within the NVIDIA Megatron-LM framework (Narayanan et al., 2021). Specifically, we run measurements on a single GPU (NVIDIA A100 80GB SXM or H100 SXM) in bfloat16 precision for Llama 2 7B and 13B. For Llama 2 70B, we run the same measurements on two GPUs of the same type with tensor parallelism. We feed the model with 2K tokens of English text and additionally generate 2K tokens in an auto-regressive manner to ensure that the model properly compresses its own generations. We limit the sequence to 4K tokens to avoid issues with context length extrapolation, as this is the maximum length observed by Llama 2 during pre-training. The reported throughput consists of the average over the last 1K tokens. To maximize the GPU utilization, we increase the batch size to the maximum that fits into memory (see Section 2.3 for details). Our implementation uses PagedAttention (Kwon et al., 2023) with a page size of $32$ tokens.
(a) Inference throughput averaged over the generation of the last 1K tokens of a 4K token sequence. On the x-axis, we show the maximum batch size that fits in memory on a single GPU (7B and 13B) or two GPUs with tensor parallelism (70B) for the vanilla LLM, and DMC 2$\times$, 4$\times$, 6$\times$, 8$\times$ models.
(b) Inference throughput–latency Pareto fronts, charted by varying the batch size with DMC 8$\times$. Up to a certain batch size, the models are memory-bound and can process additional examples with the same latency. Beyond this point, the models become increasingly compute-bound. Measurements were taken over the generation of the last 1K tokens of a 4K token sequence on an NVIDIA H100.
(c) Latency of next token generation. Solid lines denote measurements with the maximum batch size that fits on a single GPU. Dotted lines denote DMC 4$\times$ with the same batch size as the vanilla LLM.
@@ 5.2 Throughput and Latency Measurements
The compression of the KV cache with DMC allows for substantial increases in batch size and thus significant throughput gains. As shown in Figure 5(a), DMC allows for nearly linear scaling of the batch size with CR, which translates into effective increases in tokens per second compared to the original LLM. Specifically, DMC enables up to a 2.0×2.0$\times$2.0 × throughput increase for DMC 2×2$\times$2 ×, and up to a 3.9×3.9$\times$3.9 × increase for DMC 4×4$\times$4 ×. However, the scaling of throughput with CR is not linear. After reaching certain batch sizes, the models start to become compute-bound (Figure 5(b)), and the latency of inferring the next token begins to increase. Indeed, for DMC 6×6$\times$6 × and DMC 8×8$\times$8 ×, throughput starts increasing sub-linearly with CR, reaching up to $\times5$ and 7×7$\times$7 ×, respectively.
In addition, Figure 5(b) shows that increasing the batch size with DMC might trade off latency for throughput, and the Pareto front dominates the vanilla model. The different slopes of the curves on the right-hand side of Figure 5(b) stem mostly from the different hidden sizes of these models. We note that the extra memory saved with DMC could also be used to cache longer contexts.
@@
Finally, we plot how the latency changes as new tokens are generated with a fixed batch size (Figure 5(c)). When we use the largest possible batch size for either model, after generating approximately 2200 tokens, the inference time begins to scale linearly with the context size due to the increasingly dominant cost of reading the KV cache from HBM. On the other hand, if we choose to maintain the same batch size for DMC 4×4$\times$4 × as for the original LLM, the memory footprint of the KV cache is reduced and latency for longer contexts significantly improves.
@@
While we acknowledge that the behavior of LLMs at inference time depends on a multitude of factors and implementation details, our measurements in Figure 5 offer evidence that DMC can increase throughput and reduce the latency of auto-regressive generation with LLMs. We speculate that in the future, DMC might be used to grow the KV cache sub-linearly, which would provide an alternative between vanilla Transformers and State Space Models, where memory is constant (Fu et al., 2023; Gu & Dao, 2023).
5.3 Per-Head Learned Compression Ratios